40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 13:56:56 2011, kiwamu, Update, ASC, C1ASS : status update 
    Reply  Thu Sep 22 17:48:20 2011, kiwamu, Update, SUS, bad oplev on ETMY OLETMY.png
       Reply  Thu Sep 22 20:12:54 2011, kiwamu, Update, SUS, ETMY oplev whitening engaged OLETMY_WF.png
    Reply  Mon Sep 26 12:41:27 2011, kiwamu, Update, ASC, C1ASS : status update 
       Reply  Fri Sep 30 01:18:39 2011, kiwamu, Update, ASC, C1ASS : status update 
Message ID: 5518     Entry time: Thu Sep 22 13:56:56 2011     Reply to this: 5521   5543
Author: kiwamu 
Type: Update 
Category: ASC 
Subject: C1ASS : status update 

The output matrix in the C1ASS servo were coarsely readjusted and the servos seemed working.

However it is difficult to say the servo is very good or so-so,

because the ETMY suspension moves a lot and hence the cavity eigen axis moves a lot too.

 


(to do)

 + optimization of the ETMY oplevs and OSEM damping.
 + evaluation of the performance of the C1ASS with a good damping.

(Background)

 Since we have installed the new mid-HV amplifier for the PZT1 mirror (#5450) it changed the response of the PZT1 (gains from EPICS to the actual angles).
Therefore the C1ASS output matrix needs be adapted to the new PZT1 response.
 
(What I have done)
  What I was measuring was a coupling from each PZT mirror to both beam angle and beam position by looking at the output from the LOCKINs.
So this measurement eventually gives us a nicely diagonalized output matrix by inverting the coupling.
However the measurement turned out to be difficult because the ETMY moved too much.
In fact the cavity eigen axis also moves and the fluctuation was larger than the intentionally introduced beam angle/translation offsets, which are for the coupling measurement.
 
 Instead of measuring the couplings, I put some numbers into the matrix based on a guess.
Since the PZT1 HV amp became weaker than that of PZT2, the elements in the output matrix should be amplified by some number.
Right now the PZT1 amp can drive the mirror in a range of -5 -30 V with EPICS range of +/-10 counts, and for PZT2 it is about 0 -150V with EPICS range of +/-5 counts.
So the difference of the responses in unit of V/counts is about 8.5.
The PZT1 elements in the matrix were multiplied by this number and I became able to close the servos.

Quote from #5455

  + Modification of C1ASS (Kiwamu)

ELOG V3.1.3-